OpenStudy (lieutenantgeneral):

Ms. Sanchez has 26 students in her class and 14 of them are boys.What is the ratio of girls to boys? A. 6 : 13 B. 6 : 7 C. 7 : 6 D. 13 : 6

OpenStudy (solomonzelman):

26 students is the total number of srudents 14 is the number of girls there are girls and boys, so the number of boys would be b+14=26 solve the equation for b (the number of buys in the class) b+14=26 -14 -14 b= 12 "ratio of boys to girl " means boys : girls you know there are 12 boys and 14 girls, so it would be 12:14 reduce the ratio by dividing each side by 2, and you get ?

OpenStudy (anonymous):

b, with Girls to boys 12:14 is 6:7

OpenStudy (solomonzelman):

yeah!

OpenStudy (anonymous):

soloman, there's 12 girls and 14 boys lol. Ratio of girls:boys

OpenStudy (lieutenantgeneral):

thank you

OpenStudy (solomonzelman):

Sorry, it's the other way as @ECSarabia says, it's boys to girls. silly mistake, sorry :(

OpenStudy (lieutenantgeneral):

The temperature of an oven went from 320° to 280°. What was the percent decrease in temperature? A. 14.3% B. 12.5% c 1.25% D. 1.43%